please help.
Solve the following equation.
2/x + 3 = 5/x

Answer by sarah_adam(201)

2/(x+3) = 5/x
cross multipling
2x=5(x+3)
2x = 5x +15
5x - 2x = -15
3x = -15
x = -15/3 = -5